Which of the following pairs represent the same rational number? 
(-7/21) and (3/9)
 

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

We have to check the given pair represents the same rational number.
Then,
-7/21 = 3/9
-1/3 = 1/3
∵-1/3≠1/3
∴-7/21≠ 3/9
So, the given pair is not represents the same rational number.

​Thanks